A ball thrown straight up travels for 3.0s before beginning to fall. Neglecting air resistance, with what velocity was the ball thrown?
Answer by ikleyn(53763)

.
The ball started with velocity v directed vertically up, and was decelerated with the deceleration of g = 9.81 m/s^2


(the gravity acceleration at the Earth surface).


For the goals of this problem you can accept g = 10 m/s^2.


Then the equation to stop moving up is


    0 = v - gt,   or  v = g*t = 10*3 = 30 m/s.


Answer.  The initial velocity of the ball was  30 m/s.